Dog Up! Franks is looking at a new sausage system with an installed cost of $421,200. This cost will be depreciated straight-line to zero over the project's 10-year life, at the end of which the sausage system can be scrapped for $64,800. The sausage system will save the firm $129,600 per year in pretax operating costs, and the system requires an initial investment in net working capital of $30,240. Required: If the tax rate is 32 percent and the discount rate is 12 percent, what is the NPV of this project?
Answer #1

Annual depreciation = 421,200/ 10 = 42,120

Initial investment = Cost + increase in NWC

Initial investment = 421,200 + 30,240

Initial investment = 451,440

Operating cash flow from year 1 to year 10 = (Savings - depreciation)(1 - tax) + depreciation

Operating cash flow from year 1 to year 10 = (129,600 - 42,120)(1 - 0.32) + 42,120

Operating cash flow from year 1 to year 10 = 59,486.4 + 42,120

Operating cash flow from year 1 to year 10 = 101,606.4

Year 10 non operating cash flow = Market value - tax(market value - book value)

Year 10 non operating cash flow = 64,800 - 0.32(64,800 - 0)

Year 10 non operating cash flow = 64,800 - 20,736

Year 10 non operating cash flow = 44,064

NPV = Annuity * [1 - 1 / (1 + r)n] / r + FV / (1 + r)n - Initial investment

NPV = 101,606.4 * [1 - 1 / (1 + 0.12)10] / 0.12 + 44,064 / (1 + 0.12)10 - 451,440

NPV = 101,606.4 * [1 - 0.321973] / 0.12 + 14,187.4287 - 451,440

NPV = 101,606.4 * 5.650225 + 14,187.4287 - 451,440

NPV = $136,846.45